# Boson-fermion pairing and condensation in two-dimensional Bose-Fermi mixtures

### v2

@@ 4 Numerical results
In this section, we present the fully numerical results for the boson and fermion momentum distributions and for key thermodynamic quantities like the chemical potentials and condensate fraction with varying coupling strength. We also discuss the results for Tan’s contact parameter [116, 117, 118]. In all calculations we restrict to the case of equal boson and fermion masses: $m_{\rm B}=m_{\rm F}$.
@@ 4.1 Methods
For given values of $n_{\mathrm{F}},n_{\mathrm{B}},a_{\mathrm{BF}}$ and $a_{\rm BB}$ the above system is solved for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$, as follows. First, a solution is found for $\mu_{\mathrm{B}}$ from Eq. (48a) with a standard bisection method assuming a suitable ansatz for ${\mu_{\mathrm{F}}}$ and $n_{0}$. Then we consider the 2×2 system of the remaining Eqs. (48b) and (48c) as functions of $\mu_{\mathrm{F}}$ and $n_{0}$ and apply a two-dimensional secant (quasi-Newton) method whereby the approximate Jacobian matrix (which corresponds to the approximate Hessian of the total energy) is updated according to a symmetric rank 1 algorithm [119], that is a generalization of the secant method to multidimensional problems.
@@
In addition, integration in momentum space is affected by the presence of discontinuities in the integrand due to Fermi steps. Therefore, a careful analysis of the location of these discontinuities is performed in order to identify the appropriate integration intervals, for which a standard Gauss-Legendre quadrature method is adopted [120]. The largest Fermi momentum provides a natural cutoff for the momentum convolutions defining both bosonic and fermionic self-energies.
Concerning the frequency integral yielding the momentum distributions (15) and (16) we take advantage of the large frequency behavior of the self-energies (C.5-C.6) valid in the range $|\omega|\geq 100{E_{\mathrm{F}}}$, whereas for $|\omega|<100{E_{\mathrm{F}}}$ we integrate the full numerical expression (see Appendix C.2). Momentum integration in expressions (17) and (18) is carried out numerically up to a cutoff of $k_{c}=4k_{\mathrm{F}}$, after having subtracted and added a non-interacting or a Bogoliubov Green’s function respectively (see Appendix C.2), and beyond $k_{c}$ by making use of the asymptotic expressions (56-57).
@@ 4.2 Boson and fermion momentum distributions
We now discuss the numerical results for the fermionic and bosonic momentum distributions, Eqs. (15)-(16). Figure 7 shows the bosonic momentum distribution from weak to strong BF coupling for different bosonic concentrations $x$ and zero bosonic repulsion, $\eta_{\mathrm{B}}=0$. A peculiar feature common to all panels is the appearance of a peak for small but non-zero momenta, depending on the concentration and coupling. In particular, for very weak coupling the peak persists for all concentrations (full line in all panels) in contrast to higher couplings where it soon disappears. One is thus led to think that a weak-coupling mean-field effect may be at the origin of the peak, primarily due to the unpaired atoms dressing up the weakly interacting bosons, hence a sort of polaronic effect. However, for a definite identification of this peak, an analysis of the bosonic spectral weight function would be in order (this is postponed to future work).
@@
where $1/m_{\mathrm{B}}^{*}=1/m_{\mathrm{B}}+c$ and the Hugenholtz-Pines condition has been used. The coefficients $a,b,c$ are obtained by fitting the data for the bosonic self-energy at small $\omega$ and ${\mathbf{k}}$, for fixed values of $g$. Then, through a simple linear regression, one can obtain $a,b,c$ as a function of $g$, as shown in the insets of Fig. 9. One can notice that for $g\simeq-0.7$, $a=1$ and $b>0$, resulting in the following expression for the momentum distribution (16)
@@
For this reason, one needs to include a BB repulsion in order to stabilize the system against the aforementioned effects. We thus consider the case where a direct BB repulsion is turned on, causing a divergence at zero $k$ to be present at all $g$’s, as expected within Bogoliubov theory. Calculations for a relatively strong boson-boson repulsion $\eta_{\mathrm{B}}=0.1$ are reported in Fig. 10. Clearly, the main differences with respect to the case of zero repulsion occur at small momenta, where the occupation of low momenta states predicted by Bogoliubov theory dominates. Regarding the large momentum regime (not reported in Fig. 10), it is well described by expression (57) discussed below.
Concerning the Fermi momentum distribution, we first discuss its comparison with the strong-coupling expression (39) with a two-fold aim: to untangle the different contributions from atomic and molecular fermions based on what we learned in Sec. 3.2 and, at the same time, to check to what extent the approximate expression (38) for $G_{\rm F}({\mathbf{k}},\omega)$, from which Eq. (39) for ${n_{\mathrm{F}}}({\mathbf{k}})$ immediately results, provides a good approximation for $G_{\rm F}({\mathbf{k}},\omega)$. Fig. 11 displays the numerical results for ${n_{\mathrm{F}}}({\mathbf{k}})$ as symbols and the expression (39) as lines for several values of BF coupling $g$ and (a) $x=1$, (b) $x=0.5$ (c) $x=0.175$, using the same thermodynamic parameters for both calculations. The use of the same thermodynamic parameters allows for a separate check of the approximation (38) for $G_{\rm F}({\mathbf{k}},\omega)$, independently of the remaining strong-coupling approximations (46) and (47) affecting the values of the thermodynamic parameters. The agreement is remarkably good for $g\gtrsim 1$, and remains reasonable even at the lowest coupling $g=0.4$ considered in Fig. 11. This demonstrates that the fermionic Green’s function is well captured by the “BCS-like” expression (38) in an extended coupling range and not only for large values of $g$, where it is expected to hold.
We also note that the discussion on paired and unpaired fermions of Secs. 3.2 and 3.4 applies here, with the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$ representing the internal part of the molecular wave-function and the Fermi step being associated with unpaired yet correlated fermions. We notice that, when $g$ is reduced, discrepancies between the asymptotic curve and numerical data occur first for higher boson concentration, and they occur in the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$. This is related to the onset of the continuum of particle-hole excitations becoming comparable to the molecular binding energy, thus making the BCS-like expression (39) less valid.
@@ 4.3 Boson and fermion chemical potentials
Figure 13: (a): Bosonic chemical potential $\mu_{\mathrm{B}}$ (shifted by the binding energy ${\varepsilon_{0}}$), in units of $E_{\mathrm{F}}$, as a function of the BF coupling $g$, for different values of bosonic concentration $x$, and BB repulsion $\eta_{\mathrm{B}}=0$. Dotted lines on the left: weak-coupling expression (54). Dotted lines on the right: strong-coupling approximation of Sec. 3. Inset: comparison between $-\mu_{\mathrm{B}}$ and the binding energy ${\varepsilon_{0}}$ (dotted line), both in units of ${E_{\mathrm{F}}}$, as a function of $g$. (b): Corresponding fermionic chemical potential $\mu_{\mathrm{F}}$ in units of $E_{\mathrm{F}}$. Dotted lines on the left: weak-coupling expression (55) (reported for clarity only for the two extreme values of $x$). Dotted lines on the right: strong-coupling approximation of Sec. 3.
We stress that the above perturbative expressions have been tested against independent fixed-node diffusion Quantum Monte Carlo calculations in [104]. The first term in Eq. (54) arises from the standard Bogoliubov theory in 2D [94] whilst the second contribution stems from the BF interaction. One sees in Fig. 13 that in the weak-coupling regime, the numerical curves for both chemical potentials neatly approach the asymptotic limits (54,55) (dotted line on the left side). For the bosonic chemical potential, this occurs in a universal way, i.e., independently of the value of $x$, as required by Eq. (54) which does not depend on the bosonic concentration (and ${\varepsilon_{0}}$ is exponentially suppressed in this regime). In the opposite regime, both chemical potentials approach their strong-coupling benchmarks, depicted by the dotted lines on the right side of the graph, which are seen to be strongly dependent on concentration. Note that, for the boson chemical potential, Fig. 13(a) reports the quantity ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ in order to remove from the bosonic chemical potential the leading strong-coupling contribution (given by $-{\varepsilon_{0}}$). The inset avoids this subtraction and illustrates on a logarithmic scale how the bosonic chemical potential approaches the binding energy as $g$ increases, thus showing that all bosons are paired with fermions at large values of $g$.
@@
Finally, we note that we have verified that at finite $\eta_{\mathrm{B}}=0.1$ the results remain qualitatively the same, with quantitative changes affecting only ${\mu_{\mathrm{B}}}$, as also expected from the weak-coupling expression (54). The fermionic chemical potential is instead unaffected, in line with Eq. (55). On the other hand, in the strong-coupling regime, the boson-boson repulsion is essentially irrelevant due to the dominant dimer-atom repulsion. In summary, our results for the chemical potentials show that the present approach is able to recover both the pairing of bosons with fermions into composite fermions in the strong-coupling limit and the perturbative results of Ref. [104] in the opposite weak-coupling limit.
Figure 14: (a): Bosonic condensate fraction $n_{0}/n_{\mathrm{B}}$ as a function of the BF coupling parameter $g$, for different values of boson concentration $x$ and BB repulsion $\eta_{\mathrm{B}}=0$. Circles: Diagrammatic Monte Carlo results for the polaron quasiparticle residue $Z$ [99]. Squares: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ [121]. Diamonds: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ by us. (b): Square of hybridization energy scale $\Delta_{0}^{2}=2\pi{\varepsilon_{0}}n_{0}/m_{r}$ (in units of $E_{\rm F}^{2}$), as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$. Strong-coupling benchmarks from Sec. 3 are also reported as dotted lines on the right.
@@ 4.4 Boson condensate density and atom-molecule hybridization energy scale
We now discuss the results for the bosonic condensate density $n_{0}$ and the related hybridization energy scale $\Delta_{0}$. Figure 14(a) reports the value of the condensate fraction as a function of coupling (for $\eta_{\mathrm{B}}=0$, without loss of generality for this quantity). A striking feature of Fig. 14(a) is the universal behavior of $n_{0}$ with respect to concentration, in complete analogy with what found in 3D [91] and recently confirmed experimentally in [33].
@@
Such identification between the universal condensate fraction and $Z$ is challenged by the present results in 2D. One sees that, even though the condensate fraction $n_{0}/n_{\rm B}$ and the polaron quasiparticle residue $Z$ share the same qualitative behavior, significant quantitative discrepancies between the two quantities are present. We emphasize that the existence of such a discrepancy cannot be ascribed to the specific choice of the self-energy used in our calculation. One sees indeed that such a discrepancy occurs also when $Z$ is obtained with the non-self-consistent $T$-matrix approach (squares in Fig. 14(a)), which is exactly the same approach used here when carried over to the single-impurity limit. As an independent check, we have also calculated $Z$ from the derivative of the bosonic self-energy with respect to the frequency by using the same code used at finite $x$, adapted to the single-impurity limit. Our results (diamonds) fully agree with the corresponding results for $Z$ from [121], thus showing that the discrepancy between $Z$ and the condensate fraction cannot be ascribed to possible numerical errors. We thus conclude that the equivalence between the quasiparticle residue $Z$ of the Fermi polaron and the universal condensate fraction observed in 3D is only approximate, and the differences between these two quantities are amplified in 2D.
@@
As a matter of fact, and contrary to what it was argued in [91], there is no reason why the limit for $x\to 0$ of the condensate fraction and the polaron residue $Z$ should coincide. The first quantity is defined by the ratio between the number of condensed bosons $N_{0}$ and the total number of bosons $N_{\rm B}$ in the thermodynamic limit ($N_{\rm B}\to\infty$, $V\to\infty$, at fixed $n_{\rm B}$ and $x$, $V$ being the volume). It is only in this limit that a condensate fraction is well defined. So, no matter how small $x$ is, the number of bosons $N_{\rm B}$ will always scale to infinity in the thermodynamic limit, while in the polaron limit one has instead $N_{\rm B}=1$. Therefore, the occupancy of the zero-momentum state for a single impurity cannot be in general related to the ratio $N_{0}/N_{\rm B}$ in the limit $x\to 0$. It is a matter of order of limits. In the first case (condensate fraction), one first fixes $x$, takes the thermodynamic limit $V\to\infty$, and finally let $x\to 0$. In the second case (polaron $Z$), one first fixes $N_{\rm B}=1$ and then takes the limit $V\to\infty$ in such a way that $x=1/(n_{\rm F}V)\to 0$. It is thus clear that in this second case the thermodynamic limit and the limit $x\to 0$ are taken simultaneously, while in the first case the thermodynamic limit is taken first, followed by the limit $x\to 0$.
@@
A further difference with respect to the 3D case is observed in the behavior of $n_{0}$ for strong coupling. While in 3D $n_{0}$ drops identically to zero for values of $g$ larger than a critical coupling $g_{c}$, in 2D $n_{0}$ never vanishes (even though it becomes exponentially small with the coupling $g$). This in turn implies that the hybridization energy scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$ remains in general finite even at large values of $g$, except for the case $x=1$ (besides the trivial case $x=0$), as shown in Fig. 14(b). The absence of a critical coupling in 2D here obtained recalls the absence of the polaron-to-molecule transition which, as mentioned in Sec. 2.2, is a known shortcoming of the present approach when applied to the single impurity problem in 2D [97, 98]. This defect originates from an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory [97]. Such an overestimate persists also at finite $x$ (more on this point in Sec. 5). We cannot therefore exclude that the absence of a critical coupling strength in 2D is just an artifact of the present approximation, in particular at small values of $x$, which are closer to the single impurity limit.

### v3

@@ 4 Numerical results
In this section, we present the fully numerical results for the boson and fermion momentum distributions and for key thermodynamic quantities like the chemical potentials and condensate fraction with varying coupling strength. We also discuss the results for Tan’s contact parameter [118, 119, 120]. In all calculations, we consider only the case of equal boson and fermion masses: $m_{\rm B}=m_{\rm F}$.
@@ 4.1 Methods
For given values of $n_{\mathrm{F}},n_{\mathrm{B}},a_{\mathrm{BF}}$ and $a_{\rm BB}$ the above system is solved for the unknowns ${\mu_{\mathrm{F}}},{\mu_{\mathrm{B}}}$ and $n_{0}$, as follows. First, a solution is found for $\mu_{\mathrm{B}}$ from Eq. (48a) with a standard bisection method assuming a suitable ansatz for ${\mu_{\mathrm{F}}}$ and $n_{0}$. Then we consider the 2×2 system of the remaining Eqs. (48b) and (48c) as functions of $\mu_{\mathrm{F}}$ and $n_{0}$ and apply a two-dimensional secant (quasi-Newton) method whereby the approximate Jacobian matrix (which corresponds to the approximate Hessian of the total energy) is updated according to a symmetric rank 1 algorithm [121], which is a generalization of the secant method to multidimensional problems.
@@
In addition, integration in momentum space is affected by the presence of discontinuities in the integrand due to Fermi steps. Therefore, a careful analysis of the location of these discontinuities is performed to identify the appropriate integration intervals, for which a standard Gauss-Legendre quadrature method is adopted [122]. The largest Fermi momentum provides a natural cutoff for the momentum convolutions defining both bosonic and fermionic self-energies.
Concerning the frequency integral yielding the momentum distributions (15) and (16) we take advantage of the large frequency behavior of the self-energies (C.5-C.6) valid in the range $|\omega|\geq 100{E_{\mathrm{F}}}$, whereas for $|\omega|<100{E_{\mathrm{F}}}$ we integrate the full numerical expression (see Appendix C.2). Momentum integration in expressions (17) and (18) is carried out numerically up to a cutoff of $k_{c}=4k_{\mathrm{F}}$, after having subtracted and added a non-interacting or a Bogoliubov Green’s function respectively (see Appendix C.2), and beyond $k_{c}$ by making use of the asymptotic expressions (60-61).
@@ 4.2 Boson and fermion momentum distributions
We now discuss the numerical results for the fermionic and bosonic momentum distributions, Eqs. (15)-(16). Figure 7 shows the bosonic momentum distribution from weak to strong BF coupling for different bosonic concentrations $x$ and zero bosonic repulsion, $\eta_{\mathrm{B}}=0$. A peculiar feature common to all panels is the appearance of a peak for small but non-zero momenta, depending on the concentration and coupling. In particular, for very weak coupling, the peak persists for all concentrations (full line in all panels) in contrast to higher couplings where it soon disappears. One is thus led to think that a weak-coupling mean-field effect may be at the origin of the peak, primarily due to the unpaired atoms dressing up the weakly interacting bosons, hence a sort of polaronic effect. However, for a definite identification of this peak, an analysis of the bosonic spectral weight function would be in order (this is postponed to future work).
@@
where $1/m_{\mathrm{B}}^{*}=1/m_{\mathrm{B}}+c$ and the Hugenholtz-Pines condition has been used. The coefficients $a,b,c$ are obtained by fitting the data for the bosonic self-energy at small $\omega$ and ${\mathbf{k}}$, for fixed values of $g$. Then, through a simple linear regression, $a,b,c$ can be obtained as a function of $g$, as shown in the insets of Fig. 9. One can notice that for $g\simeq-0.7$, $a=1$ and $b>0$, resulting in the following expression for the momentum distribution (16)
@@
For this reason, one needs to include a BB repulsion in to stabilize the system against the aforementioned effects. We thus consider the case where a direct BB repulsion is turned on, causing a divergence at zero $k$ to be present at all $g$’s, as expected within Bogoliubov theory. Calculations for a relatively strong boson-boson repulsion $\eta_{\mathrm{B}}=0.1$ are reported in Fig. 10. Clearly, the main differences with respect to the case of zero repulsion occur at small momenta, where the occupation of low momenta states predicted by Bogoliubov theory dominates. Regarding the large momentum regime (not reported in Fig. 10), it is well described by expression (61) discussed below.
Concerning the Fermi momentum distribution, we first discuss its comparison with the strong-coupling expression (39) with a two-fold aim: to untangle the different contributions from atomic and molecular fermions based on what we learned in Sec. 3.2 and, at the same time, to check to what extent the approximate expression (38) for $G_{\rm F}({\mathbf{k}},\omega)$, from which Eq. (39) for ${n_{\mathrm{F}}}({\mathbf{k}})$ immediately results, provides a good approximation for $G_{\rm F}({\mathbf{k}},\omega)$. Fig. 11 shows the numerical results for ${n_{\mathrm{F}}}({\mathbf{k}})$ as symbols and the expression (39) as lines for several values of BF coupling $g$ and (a) $x=1$, (b) $x=0.5$ (c) $x=0.175$, using the same thermodynamic parameters for both calculations. The use of the same thermodynamic parameters allows for a separate check of the approximation (38) for $G_{\rm F}({\mathbf{k}},\omega)$, independently of the remaining strong-coupling approximations (46) and (47) affecting the values of the thermodynamic parameters. The agreement is remarkably good for $g\gtrsim 1$, and remains reasonable even at the lowest coupling $g=0.4$ considered in Fig. 11. This demonstrates that the fermionic Green’s function is well captured by the “BCS-like” expression (38) in an extended coupling range and not only for large values of $g$, where it is expected to hold.
We also note that the discussion on paired and unpaired fermions of Secs. 3.2 and 3.4 applies here, with the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$ representing the internal part of the molecular wave-function and the Fermi step being associated with unpaired yet correlated fermions. We notice that, when $g$ is reduced, discrepancies between the asymptotic curve and the numerical data occur first for the higher boson concentration, and they occur in the tail of ${n_{\mathrm{F}}}({\mathbf{k}})$. This is related to the onset of the continuum of particle-hole excitations becoming comparable to the molecular binding energy, thus making the BCS-like expression (39) less valid.
@@ 4.3 Boson and fermion chemical potentials
Figure 13: (a): Bosonic chemical potential $\mu_{\mathrm{B}}$ (shifted by the binding energy ${\varepsilon_{0}}$), in units of $E_{\mathrm{F}}$, as a function of the BF coupling $g$, for different values of bosonic concentration $x$, and BB repulsion $\eta_{\mathrm{B}}=0$. Dotted lines on the left: weak-coupling expression (54). Dotted lines on the right: strong-coupling approximation of Sec. 3. Inset: comparison between $-\mu_{\mathrm{B}}$ and the binding energy ${\varepsilon_{0}}$ (dotted line), both in units of ${E_{\mathrm{F}}}$, as a function of $g$. (b): Corresponding fermionic chemical potential $\mu_{\mathrm{F}}$ in units of $E_{\mathrm{F}}$. Dotted lines on the left: weak-coupling expression (55) (reported for clarity only for $x=1$ and $x=0.175$). Dotted lines on the right: strong-coupling approximation of Sec. 3.
We stress that the above perturbative expressions have been tested against independent fixed-node diffusion Quantum Monte Carlo calculations in [104]. The first term in Eq. (54) arises from the standard Bogoliubov theory in 2D [94], while the second contribution stems from the BF interaction. One sees in Fig. 13 that in the weak-coupling regime, the numerical curves for both chemical potentials neatly approach the asymptotic limits (54,55) (dotted line on the left side). For the bosonic chemical potential, this occurs in a universal way, i.e., independently of the value of $x$, as required by Eq. (54) which does not depend on the bosonic concentration (and ${\varepsilon_{0}}$ is exponentially suppressed in this regime). In the opposite regime, both chemical potentials approach their strong-coupling benchmarks, depicted by the dotted lines on the right side of the graph, which are seen to be strongly dependent on concentration. Note that, for the boson chemical potential, Fig. 13(a) reports the quantity ${\mu_{\mathrm{B}}}+{\varepsilon_{0}}$ to remove from the bosonic chemical potential the leading strong-coupling contribution (given by $-{\varepsilon_{0}}$). The inset avoids this subtraction and illustrates on a logarithmic scale how the bosonic chemical potential approaches the binding energy as $g$ increases, thus showing that all bosons are paired with fermions at large values of $g$.
@@
Figure 14: (a): Bosonic chemical potential $\mu_{\mathrm{B}}$ (shifted by the binding energy ${\varepsilon_{0}}$), in units of $E_{\mathrm{F}}$, as a function of the BF coupling $g$, for different values of bosonic concentration $x$, and BB repulsion $\eta_{\mathrm{B}}=0.1$. Dotted lines on the left: weak-coupling expression (54) (reported for clarity only for $x=1$ and $x=0.175$) . Dotted lines on the right: strong-coupling approximation of Sec. 3. (b): Corresponding fermionic chemical potential $\mu_{\mathrm{F}}$ in units of $E_{\mathrm{F}}$. Dotted lines on the left: weak-coupling expression (55) (reported for clarity only for $x=1$ and $x=0.175$). Dotted lines on the right: strong-coupling approximation of Sec. 3.
Finally, Fig.14 shows that at finite $\eta_{\mathrm{B}}=0.1$ the results remain qualitatively the same, with quantitative changes that affect only ${\mu_{\mathrm{B}}}$, as also expected from the weak-coupling expression (54). The fermionic chemical potential is instead unaffected, in line with Eq. (55). On the other hand, in the strong-coupling regime, the boson-boson repulsion is essentially irrelevant due to the dominant dimer-atom repulsion. In summary, our results for the chemical potentials show that the present approach is able to recover both the pairing of bosons with fermions into composite fermions in the strong-coupling limit and the perturbative results of Ref. [104] in the opposite weak-coupling limit.
@@ 4.4 Boson condensate density and atom-molecule hybridization energy scale
Figure 15: (a): Bosonic condensate fraction $n_{0}/n_{\mathrm{B}}$ as a function of the BF coupling parameter $g$, for different values of boson concentration $x$ and BB repulsion $\eta_{\mathrm{B}}=0$. Circles: Diagrammatic Monte Carlo results for the polaron quasiparticle residue $Z$ [99]. Squares: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ [123]. Diamonds: non-self-consistent $T$-matrix approximation results for the polaron quasiparticle residue $Z$ by us. (b): Square of hybridization energy scale $\Delta_{0}^{2}=2\pi{\varepsilon_{0}}n_{0}/m_{r}$ (in units of $E_{\rm F}^{2}$), as a function of the BF coupling $g$, for different values of the bosonic concentration $x$ and boson-boson repulsion $\eta_{\mathrm{B}}=0$. Strong-coupling benchmarks from Sec. 3 are also reported as dotted lines on the right. (c): Bosonic condensate fraction $n_{0}/n_{\mathrm{B}}$ as a function of the BF coupling parameter $g$, for different values of boson concentration $x$ and BB repulsion $\eta_{\mathrm{B}}=0.1$.
We now discuss the results for the bosonic condensate density $n_{0}$ and the related hybridization energy scale $\Delta_{0}$. Figure 15 reports the value of the condensate fraction as a function of coupling (for $\eta_{\mathrm{B}}=0$ in panel (a) and $\eta_{\mathrm{B}}=0.1$ in panel (c)). A striking feature of Fig. 15 is the nearly universal behavior of $n_{0}$ with respect to concentration, in complete analogy to what was found in 3D [91] and recently confirmed experimentally in [33].
@@
Such identification between the universal condensate fraction and $Z$ is challenged by the present results in 2D. One sees that, even though the condensate fraction $n_{0}/n_{\rm B}$ and the polaron quasiparticle residue $Z$ share the same qualitative behavior, significant quantitative discrepancies between the two quantities are present. We emphasize that the existence of such a discrepancy cannot be ascribed to the specific choice of the self-energy used in our calculation. One sees indeed that such a discrepancy occurs also when $Z$ is obtained with the non-self-consistent $T$-matrix approach (squares in Fig. 15(a)), which is exactly the same approach used here when carried over to the single-impurity limit. As an independent check, we have also calculated $Z$ from the derivative of the bosonic self-energy with respect to the frequency by using the same code used at finite $x$, adapted to the single-impurity limit. Our results (diamonds) fully agree with the corresponding results for $Z$ from [123], thus showing that the discrepancy between $Z$ and the condensate fraction cannot be ascribed to possible numerical errors. We therefore conclude that the equivalence between the quasiparticle residue $Z$ of the Fermi polaron and the universal condensate fraction observed in 3D is only approximate, and the differences between these two quantities are amplified in 2D.
In this respect, reconsidering the 3D case and focusing on unitarity, by comparing the data from [91] for $n_{0}/n_{\rm B}$ at the lowest concentration ($x=0.175$) with the polaron residue obtained with the non-self-consistent T-matrix approach [124] ($Z^{\rm TMA}$) or with the diagrammatic Monte-Carlo method [124] ($Z^{\rm dMC}$), one has $n_{0}/n_{\rm B}=0.74$, $Z_{\rm pol}^{\rm TMA}=0.80$, $Z^{\rm dMC}=0.76$ for $m_{\rm B}/m_{\rm F}=0.575$; $n_{0}/n_{\rm B}=0.73$, $Z^{\rm TMA}=0.78$, $Z^{\rm dMC}=0.76$ for $m_{\rm B}/m_{\rm F}=1$; $n_{0}/n_{\rm B}=0.60$, $Z^{\rm TMA}=0.67$, $Z^{\rm dMC}=0.65$ for $m_{\rm B}/m_{\rm F}=5$. One sees that $n_{0}/n_{\rm B}$ and $Z^{\rm TMA}$ differ by approximately 6% for equal masses, and the difference increases by changing the mass ratio, exceeding 10% for $m_{\rm B}/m_{\rm F}=5$. (The difference between $n_{0}/n_{\rm B}$ and $Z^{\rm dMC}$ is slightly smaller, but the comparison with $Z^{\rm TMA}$ is more meaningful, since $Z^{\rm TMA}$ and $n_{0}/n_{\rm B}$ are calculated within the same approximation.) This further indicates that the equivalence observed in 3D between the polaron residue and $n_{0}/n_{\rm B}$ is only approximate; the present calculation indicates that this approximate degeneracy is lifted in 2D.
@@
In fact, and contrary to what was argued in [91], there is no reason why the limit for $x\to 0$ of the condensate fraction should coincide with the polaron residue $Z$. The first quantity is defined by the ratio between the number of condensed bosons $N_{0}$ and the total number of bosons $N_{\rm B}$ in the thermodynamic limit ($N_{\rm B}\to\infty$, $V\to\infty$, at fixed $n_{\rm B}$ and $x$, $V$ being the volume). Only in this limit is a condensate fraction well defined. So, no matter how small $x$ is, the number of bosons $N_{\rm B}$ will always scale to infinity in the thermodynamic limit, while in the polaron limit one has instead $N_{\rm B}=1$. Therefore, the occupancy of the zero-momentum state for a single impurity cannot be in general related to the ratio $N_{0}/N_{\rm B}$ in the limit $x\to 0$. It is a matter of the order of limits. In the first case (condensate fraction), one first fixes $x$, takes the thermodynamic limit $V\to\infty$, and finally lets $x\to 0$. In the second case (polaron $Z$), one first fixes $N_{\rm B}=1$ and then takes the limit $V\to\infty$ in such a way that $x=1/(n_{\rm F}V)\to 0$. Hence, in this second case the thermodynamic limit and the limit $x\to 0$ are taken simultaneously, while in the first case the thermodynamic limit is taken first, followed by the limit $x\to 0$. It is thus clear that the two limits are different:
$\lim_{x\to 0}\lim_{V\to\infty}\frac{n_{\rm B}(k=0;N_{\rm B}=xN_{\rm F};N_{\rm F}/V={\rm const})}{N_{\rm B}}\neq\lim_{V\to\infty}\frac{n_{\rm B}(k=0;N_{\rm B}=1;N_{\rm F}/V={\rm const})}{1}.$ (56)
Figure 16: Imaginary part of the bosonic self-energy, in units of $E_{\rm F}$, as a function of the frequency $\omega$ along the imaginary axis for different values of boson concentration $x$ at BF coupling $g=0$ and BB repulsion $\eta_{\rm B}=0$. The case $x=0$ corresponds to the calculation in the polaron limit of a single impurity in a Fermi bath.
A interesting related question is how the polaron residue $Z$ is obtained from the limiting behavior of the bosonic self-energy when $x\to 0$. One has [125, 117]
$Z=\lim_{\omega^{\prime}\to 0}\frac{1}{\left|1-\frac{\partial{\rm Re}\Sigma^{\rm R}_{\rm p}(0,\omega^{\prime})}{\partial\omega^{\prime}}\right|}=\lim_{\omega\to 0^{+}}\frac{1}{\left|1-\frac{\partial{\rm Im}\Sigma_{\rm p}(0,\omega)}{\partial\omega}\right|}.$ (57)
Here, $\Sigma_{\rm p}^{\rm R}(0,\omega^{\prime})$ is the retarded polaron self-energy calculated at zero momentum and at the real frequency $\omega^{\prime}$, while $\Sigma_{\rm p}(0,\omega)$ is its analytic extension into the upper complex plane calculated at the imaginary frequency $i\omega$. For a given BF coupling strength, one has $\Sigma_{\rm p}(0,\omega)=\lim_{x\to 0}\Sigma_{\rm B}(0,\omega;x)$, where $\Sigma_{\rm B}(0,\omega;x)$ is the boson self-energy at finite $x$ calculated at the imaginary frequency $i\omega$, for $\eta_{\rm B}=0$. Hence,
$Z=\lim_{\omega\to 0^{+}}\lim_{x\to 0}\frac{1}{\left|1-\frac{\partial{\rm Im}\Sigma_{\rm B}(0,\omega;x)}{\partial\omega}\right|}.$ (58)
One may also consider, at finite $x$, the quasiparticle residue of the bosons at vanishig momentum and frequency:
$Z(x)=\lim_{\omega\to 0^{+}}\frac{1}{\left|1-\frac{\partial{\rm Im}\Sigma_{\rm B}(0,\omega;x)}{\partial\omega}\right|}.$ (59)
It is evident that in general $Z$ might be different from $\lim_{x\to 0}Z(x)$ due to the different order in the two limits involved. That this is indeed the case, it is suggested by the results of Sec. 4.2 (in particular the discussion of Fig. 9), which show that $Z(x)\gg 1$ around $g=-0.7$, against a corresponding value of $Z\simeq 0.9$.
To further analyze this behavior at a more generic BF coupling, Fig. 16 reports ${\rm Im}\Sigma_{\rm B}(0,\omega)$ at $g=0$ for three different concentrations approaching the limit $x\to 0$, together with its expected limiting value for $x=0$, ${\rm Im}\Sigma_{\rm p}(0,\omega)$. One sees that ${\rm Im}\Sigma_{\rm B}(0,\omega)$ tends indeed to ${\rm Im}\Sigma_{\rm p}(0,\omega)$. However, the slope of ${\rm Im}\Sigma_{\rm B}(0,\omega)$ in $\omega=0$ is nearly independent of $x$ and, for $x\to 0$, tends to a value that is completely different from the slope of ${\rm Im}\Sigma_{\rm p}(0,\omega)$ in $\omega=0$. In this way, one has $\lim_{x\to 0}Z(x)=0.93$, against a value of $Z=0.54$.
@@
Returning to Fig. 15, another difference is observed with respect to the 3D case in the behavior of $n_{0}$ for strong coupling. While in 3D $n_{0}$ drops identically to zero for values of $g$ larger than a critical coupling $g_{c}$, in 2D $n_{0}$ never vanishes (even though it becomes exponentially small with the coupling $g$). This in turn implies that the hybridization energy scale $\Delta_{0}\propto(n_{0}{\varepsilon_{0}})^{1/2}$ remains in general finite even at large values of $g$, except for the case $x=1$ (besides the trivial case $x=0$), as shown in Fig. 15(b). The absence of a critical coupling in 2D here obtained recalls the absence of the polaron-to-molecule transition, which, as mentioned in Sec. 2.2, is a known shortcoming of the present approach when applied to the single impurity problem in 2D [97, 98]. This defect originates from an overestimate of the atom-molecule repulsion in the strong-coupling limit of the theory [97]. This overestimate persists also at finite $x$ (more on this point in Sec. 5). We cannot therefore exclude that the absence of a critical coupling strength in 2D is just an artifact of the present approximation, in particular at small values of $x$, which are closer to the single impurity limit.